on August 15 Sheffield Company discounted at sunshine bank a $8,500 (maturity value), 130- day note dated May 13. Sunshine’s discount rate was 10%. (Use days in a year table) what proceeds did Sheffield Company receive? ( use 360 days a year. Do not round immediate calculations)
Answer #1

Difference between dates August 15 and May 13 = 08/15 - 05/13 = 94 days

Days left in the Note = 130 days - 94 days = 36 days

The Proceeds the company will receive = Maturity Value - Maturity Value * Rate of interest * 36 / 360

The Proceeds the company will receive = 8500 - 8500 * 0.10 * 36 / 360

The Proceeds the company will receive = $8500 - $85

The Proceeds the company will receive = $8415